OpenStudy (anonymous):

Simplify (6a2 + 5b)2.

OpenStudy (zepp):

\((6a^2+5b)^2\)?

OpenStudy (zepp):

Use the algebraic property: \((a+b)^2=a^2+2ab+b^2\)

OpenStudy (anonymous):

Zepp you can handle this.

OpenStudy (zepp):

\((6a^2+5b)^2=(6a^2)^2+2(6a^2+5b)+(5b)^2)\)

OpenStudy (zepp):

Simplify that by using the power or a power rule: \(\large (a^b)^c=a^(bc)\)

OpenStudy (zepp):

\(a^{bc}\), sorry

OpenStudy (zepp):

\((6a^2)^2+2(6a^2+5b)+(5b)^2=36a^4+12a^2+10b+25b^2\)

OpenStudy (anonymous):

A:36a2 + 60a4b + 25b2 C:36a4 + 60a4b2 + 25b2 B:36a4 + 30a4b + 25b2 D:36a4 + 60a2b + 25b2

OpenStudy (zepp):

Oups, the middle term should be \(2(6a^2*5b)=60a^2b\)

OpenStudy (zepp):

So D)